OpenStudy (ashy98):

The price of products may increase due to inflation and decrease due to depreciation. Derek is studying the change in the price of two products, A and B, over time. The price f(x), in dollars, of product A after x years is represented by the function below. f(x) = 0.69(1.03)x Part A: Is the price of product A increasing or decreasing and by what percentage per year? Justify your answer. Part B: The table below shows the price f(t), in dollars, of product B after t years: t (number of years) 1 2 3 4 f(t) (price in dollars) 10,100 10,201 10,303.01 10,406.04

OpenStudy (ashy98):

Which product recorded a greater percentage change in price over the previous year?

OpenStudy (anonymous):

is the function .69*1.03*x?

OpenStudy (ashy98):

12500(0.82)=10250 so it is decreasing by 2250 its part A The highest percent would be it because for product a its 2550 then you have to plug it in so for product A=12500(0.82) times year 2 would be 12500(0.82)^2 would be 12500(1.62)=20,250 so therefore it would be increasing. right??!!!?!?!

OpenStudy (anonymous):

for part A its decreasing, but by .69*1.03=.71

OpenStudy (anonymous):

product b tends to increase

OpenStudy (anonymous):

yes you are right

OpenStudy (ashy98):

okay thanks (:

OpenStudy (hahawowoh):

what was part A?